You have $10,000 that you plan to invest in a compound-interest bearing interest. Your investment agent advises you that you can invest the $10,000 at 8% compounded quarterly for three years or you can invest it at 8 ¼ % compounded annually for three years. Which investment should you choose to receive the most interest? Show your work for calculating the interest for both options.
Answer by rfer(16322)

A=10000(1+0.08/4)^3*4
A=10000(1.02)^12
A=10000*1.26824
A=12,682.40
---------------
A=10000(1.0825)^3
A=10000*1.26848
A=12,684.80